A population consists of the four numbers 1, 2, and 4. List all the possible samples of size n=2 which can be drawn replacement from the population. Find the following



a.) Population means


b.) Population Variance


c.) Population Standard Deviation


d.) Mean of the sampling distribution of sample means


e.) Variance of the sampling distribution of sample means


f.) Standard deviation of the sampling distribution of sample means

1
Expert's answer
2022-03-01T18:39:39-0500

We have population values "1,2,4,5" population size "N=4," and sample size "n=2." Thus, the number of possible samples which can be drawn with replacement is "4^2=16."

a)



"mean=\mu=\dfrac{1+2+4+5}{4}=3"

b)


"Variance=\sigma^2=\dfrac{1}{4}((1-3)^2+(2-3)^2"




"+(4-3)^2+(5-3)^2)=2.5"

c)



"\sigma=\sqrt{\sigma^2}\approx1.581139"

d)


"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 No & Sample & Mean \\ \hline
 1 & (1,1) & 1 \\
 \hdashline
 2 & (1,2) & 1.5 \\
 \hdashline
3 & (1,4) & 2.5 \\
 \hdashline
4 & (1,5) & 3 \\
 \hdashline
5 & (2,1) & 1.5 \\
 \hdashline
6 & (2,2) & 2 \\
 \hdashline
7 & (2,4) & 3 \\
 \hdashline
8 & (2,5) & 3.5 \\
 \hdashline
9 & (4,1) & 2.5 \\
 \hdashline
10 & (4,2) & 3 \\
 \hdashline
11 & (4,4) & 4 \\
 \hdashline
12 & (4,5) & 4.5 \\
 \hdashline
13 & (5,1) & 3 \\
 \hdashline
14 & (5,2) & 3.5 \\
 \hdashline
15 & (5,4) &4.5 \\
 \hdashline
16 & (5,5) & 5 \\
 \hdashline
\end{array}"


The sampling distribution of the sample mean "\bar{x}" and its mean and standard deviation are:


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:}
 \bar{x} & f & f(\bar{x})& \bar{x}f(\bar{x})& \bar{x}^2f(\bar{x}) \\ \hline
 1& 1 & 1/16 & 1/16 & 2/32 \\
 \hdashline
 1.5 & 2 & 2/16 & 3/16 & 9/32 \\
 \hdashline
 2 & 1 & 1/16 & 2/16 & 8/32 \\
 \hdashline
 2.5 & 2 & 2/16 & 5/16 & 25/32 \\
 \hdashline
 3 & 4 & 4/16 & 12/16 & 72/32 \\
 \hdashline
 3.5 & 2 & 2/16 & 7/16 & 49/32 \\
 \hdashline
 4 & 1 & 1/16 & 4/16 & 32/32 \\
 \hdashline
 4.5 & 2 & 2/16 & 9/16 & 81/32 \\
 \hdashline
 5 & 1 & 1/16 & 5/16 & 50/32 \\
 \hdashline
 Total & 16 & 1 & 3 & 41/4 \\
 \hdashline
\end{array}"




"\mu_{\bar{X}}=E(\bar{X})=\sum\bar{x}f(\bar{x})=5"

e)


"Var(\bar{X})=\sigma_{\bar{X}}^2=\sum\bar{x}^2f(\bar{x})-(\sum\bar{x}f(\bar{x}))^2""=\dfrac{41}{4}-3^2=\dfrac{5}{4}"

f)



"\sigma_{\bar{X}}=\sqrt{\sigma_{\bar{X}}^2}=\sqrt{\dfrac{5}{4}}=\dfrac{\sqrt{5}}{2}\approx1.118034"



Check


"\mu_{\bar{X}}=E(\bar{X})=5=\mu"




"Var(\bar{X})=\dfrac{5}{4}=\dfrac{\sigma^2}{n}=\dfrac{2.5}{2}"